Telescope and AGw Commissioning at LFBG with IRTC - 15-16 April 2009 MST

Observers: D.Thompson, MPedani (@LBT)
Telescope Operator: W.Wack
Telescope Support: JLittle
Instrument Support: JMorris
SW Support:

Summary

  • Three (so far!) complete failures of the chillers. LUCIFER cuts off CCC power due to over-temperatures. See
  • Lost entire night to high winds (up to 88 mph this afternoon) and humidity.
  • John most likely found the source of the high residuals - truncation on conversion to radians somewhere in the path between IRC and PCS. Once fixed, we can test it closed dome.

Details

We have had a 10C drop in temperature since this afternoon. M1 is -0.8C and outside is -9.6C.

-- DavidThompson - 15 Apr 2009
Topic revision: r3 - 16 Apr 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback